Operations Report – January 25th

Operations Report

Subject Line: Crew 187 Operations Report 25JAN2018

Crew 187 Operations Report 25012018

SOL: 13

Name of person filing report: Luis Jose Antonio Diaz Lopez

Non-nominal systems: None.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15.8h

Generator turned off, charging battery at 10h:40m

Generator turned on at 19h:43m

Solar— SOC

@ 10:10: 68%

@ 19:43: 37%

Diesel: 96%

Propane: 89%

Ethanol Free Gasoline (5 Gallon containers for ATV): 10 Gallons

Water (trailer): 0 Gallons

Water (static): 140 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01302996 Gallons

Toilet tank emptied: No

ATVs Used: HONDA, 350(1), 350(2), 350(3)

Oil Added? No

ATV Fuel Used: 6%

# Hours the ATVs were used today: 01h 35min

Notes on ATVs: None

Deimos rover used: Taked by the director to testing due the recently battery problems.

Hours: 102.3 h

Beginning charge: 100%

Ending charge: –

Currently charging: Maybe

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 14.5h

Beginning charge: 100%

Ending charge: –

Currently charging: No

Opportunity rover used: Yes

Hours: 11.7h

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: 5.9h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Opportunity have more hours than the other rovers because it was left turned on by error during the before EVA. We got 10 gallons of gasoline as backup for ATV’s.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of EVA operations: All nominal

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal.

Summary of RAM operations: Not Operational. Some generator supplies were delivered into the RAM.

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: None

Sol Summary – January 25th

Crew 187 Sol 13 Summary Report 25JAN2018

Sol 13

Summary Title: Breaking the habits

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Hi Earth,

How is everything going on there?

Last night we went to see the starry sky of Mars; however, the cloudy sky did not allow us to observe the beautiful Martian sky. There are only a few days left on Mars, if it depends on us, we would stay on Mars for much longer. The fatigue is felt in the Martians; on Mars, the hours pass faster than Earth. We slept late talking, coordinating things, moving forward in our projects and so on. Today we woke up at different times, but at 9:30 we were all ready to take breakfasts. One hour later, the Martian explorers, Atila, Luis, Tania and David started their adventure. Our explorers arrived at Mountain Goat Trail when they realized that they were in the wrong way to Candor Chasma. Then, they came back almost to Galileo Road 104… OMG!… again! The Martian explorers, couldn’t find the right way, as yesterday, the third one is the defeated. Therefore, at 12:03, they arrived at the right point and walked for 100 minutes until they came back to the rovers.

Meanwhile, in the Hab, Danton, Oscar and I got ready for the gala dinner. Of course, the chef in charge of the kitchen was our Martian chef Oscar. Danton and I were in charge of the cleaning. When the Martian explorers arrived at the Hab, we ate lunch together. Later, David helped me taking pictures for my project. At 15:55, Tania, Luis and I were in the airlock ready for the Extraordinary EVA, we arrived until the intersection between White Moon and Factory Butte View before we came back to our Martian Home.

Currently, we are doing the reports for Capcom and getting ready for our gala dinner. Today our special guest is Dr. Ruppert, we hope she enjoys her visit in our home.

In 2 hours, we are planning to use the Celestron Telescope to see the Martian Sky. Unfortunately, yesterday we couldn’t see anything at night, the sky was cloudy.

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Working in our Martian projects.

Anomalies in work:


Weather: Temperature: 2°C, Fairweather, Winds speed: NNE 22.5 MPH, Humidity 47%, Barometer 29.81 in

Crew Physical Status: Healthy

Today’s EVA:

Location: Candor Chasma

Duration: From 10:35 until 14:10

Purpose: Testing of long-range EVA protocols and testing of temperature regulation vest.

Today’s Extraordinary EVA:

Location: Intersection between White Moon and Factory Butte View

Duration: From 15:55 until 17:00

Purpose: Usage of rovers for battery discharge.

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report and Journalist Report.

Support Requested:


Greenhab Report – January 25th

GreenHab Report

Hernán David Mateus Jiménez


Environmental control:


Ambient with door opening

Shade cloth on

Working Hour: 18:00
Inside temp at working hour: 14° C
Outside temp during working hours: 5° C
Inside temperature high: 30° C
Inside temperature low: 14° C
Inside humidity: 88 %RH

Inside humidity high: 88 %RH
Inside humidity low: 31 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: none

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 19:15

Changes to research plants: The plans of quinoa over the lysimeter are withered

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: none

Copyright © The Mars Society. All rights reserved. | Main Site